Find the smallest positive number t such that
e^(sin(t))=5/7

Answer by stanbon(75887)

Find the smallest positive number t such that
e^(sin(t))=5/7
-----
Take the natural log of both sides to get:
sin(t) = ln(5/7)
---
sin(t) = -0.3365
---
Take the inverse sin of both sides to get:
t = 200.07 degrees or 3.4848 radians
==================
Cheers,
Stan H.
==================